# Global Balance and Systemic Risk in Financial Correlation Networks

### v1

@@ Abstract
We show that the global balance index of financial correlation networks can be used as a systemic risk measure. We define the global balance of a network starting from a diffusive process that describes how the information spreads across nodes in a network, providing an alternative derivation to the usual combinatorial one. The steady state of this process is the solution of a linear system governed by the exponential of the replication matrix of the process. We provide a bridge between the numerical stability of this linear system, measured by the condition number in an opportune norm, and the structural predictability of the underlying signed network. The link between the condition number and related systemic risk measures, such as the market rank indicators, allows the global balance index to be interpreted as a new systemic risk measure. A comprehensive empirical application to real financial data finally confirms that the global balance index of the financial correlation network represents a valuable and effective systemic risk indicator.
@@
Networks, Signed networks, Global balance, Structural stability, Systemic risk measures
@@ 1 Introduction
Starting from the 2008 global financial crisis, frequent and severe crises affected the economic system. This fact renewed the interest of practitioners and academicians for systemic risk measures. A unique definition of systemic risk is missing in the literature, probably for its complexity and for the multitude of potential triggering events: financial crises, sovereign debt crises, pandemics, wars and so on. Therefore, systemic risk has been studied from many different and alternative point of views: for example, [1] proposes a macroeconomic approach; [2] and [3] applies a network connectedness analysis; [4], [5], [6] and [7], among others, focus on conditional value at risk as a measure of systemic risk; and [8] investigates the inter-bank financial network as a source of systemic risk. An extensive review of the literature on the topic is behind the scope of the present research and for more details we refer, for instance, to [9], [10] and [11].
@@
In the literature on financial networks, correlations are sometimes transformed into distances. The usual methodology is based on the nonlinear transformation $d(i,j)=\sqrt{2(1-c_{ij})}$, introduced in [15], where $c_{ij}$ is the Pearson correlation between $i$ and $j$. Other distances have been proposed to overcome some drawbacks of the previous one, such as $d(i,j)=-\log c_{ij}^{2}$ in [16]. While these transformations permit to construct weighted undirected networks and apply classical network theory tools, they fail to exploit the information contained in the correlation sign.
@@
The theory of signed networks [37] and their balance can provide interesting insight into the subtle relationship between predictability and overall systemic risk associated with groups of stocks or the entire network. A key insight, due to Harary, is that the level of structural balance impacts also on the performance of an investment portfolio. In fact, to the best of our knowledge, the first paper that or establishes a connection between signed networks and risk in financial networks is the one by Harary et al. [38]. In this paper, the authors address the question of how to design an effective hedging strategy for portfolio selection. They analyze the network of correlations between securities and conclude that a hedging strategy is effective, and thus the portfolio less risky, if there is at least one negative link and if the signed network is structurally balanced. However, the authors do not deal with systemic risk measures in the proper sense and do not study the relationship between these measures and structural balance indices. Indeed, the global balance index has never been interpreted as an indicator of systemic risk.
@@
In this research, we show how to move from these structural indicators for signed networks to systemic risk measures as usually interpreted in the literature. The logical path of the paper can be described as follows. We start from a non-conservative diffusive process on networks governed by the adjacency matrix, interpreted as a replication matrix of the information transmitted by the nodes. We interpret the local version of the structural balance as a measure of the amount of information that leaves one node and returns to the same node when the diffusion process has reached a steady state. From the local measure we move to the global balance index, which is known and widely used in the literature to quantify structural balance in signed networks. Then, we show that this global indicator is naturally related to a ratio of the condition numbers of the two linear systems defining the steady state of the diffusive process on the signed network and on the corresponding unsigned network, that is the network with all positive edges. Thus, a natural bridge is created between the structural balance in correlation networks and a measure of the sensitivity of the network to an external perturbation of the state of the nodes. The link to known algebraic measures of conditioning provides the further connection to systemic risk measures. The condition number of the correlation matrix and some suitable generalizations called market rank indicators (MRI) have already been proposed in the literature as systemic risk measures (see [39]). We then compare the global balance of the signed correlation network with these systemic risk measures, which are part of the broader class of so-called proper measures of connectedness. This comparison leads us to explore the ability of the global balance to discriminate systemic events. Finally, we look at how such a global indicator performs as an effective measure of systemic risk, testing its performance on real financial data.
@@
The paper is structured as follows: In Section 2 we recall the mathematical definitions. In Section 3 we show how the structural balance can be obtained from a discrete diffusive process on the network. In Section 4 we outline a conceptual path from the balance indices to systemic risk measures. In Section 5 we perform an empirical analysis to prove the effectiveness of the balance index as systemic risk measure. In Section 6 we draw conclusions and outlines for future research.
@@ 2 Preliminaries
We present first some preliminary notions on signed graphs and correlation networks.
@@ 2.1 Signed networks and structural balance
An undirected signed network is represented by a graph $G=(V,E,\mathbf{W})$, where $V$ is the set of $N$ nodes, and $E$ is the set of undirected edges. The weighted signed adjacency matrix $\mathbf{W}=[w_{ij}]$ of order $N\times N$, with $w_{ij}\in\mathbb{R}$, encodes all the adjacency relationships within the nodes. Specifically, $w_{ij}\neq 0$ if there is an edge between nodes $i$ and $j$, $w_{ij}=0$ otherwise. Edge weights are positive or negative depending on the nature of the relationship between the pair of nodes. It may relate to activating or inhibiting functions in biological systems ([40]), trusting or distrusting links in social or political networks ([41]), cooperative or antagonistic relationships in economic settings ([42]), correlated and anticorrelated behavior in the time series of asset returns in the financial context ([43]). A binary adjacency matrix $\mathbf{A}=[a_{ij}]$ can be associated with the signed network $G$ by setting all non-null weights equal to $1$ or $-1$ according to the sign of the entries $a_{ij}$. In the paper we will also need to refer to the underlying unsigned network, that is the graph obtained from $G$ by neglecting the edges sign. The matrices associated with this network are $|{\bf W}|$ and $|{\bf A}|$. The strength of a node $i$, in the signed network, is then defined as $s_{i}=\sum_{j}|w_{ij}|$ and the degree as $d_{i}=\sum_{j}|a_{ij}|$.
A signed graph $G$ is structurally balanced if there are no negative cycles, that is, cycles with an odd number of negative edges (see [17, 18]). A balanced network is characterized by the balance theorem as a graph containing a bipartition of the vertex set $V$ such that every edge between the two subsets is negative, while every edge within each subset is positive (see [17]).
@@ 2.2 Correlation networks from asset returns
Let us consider the time series of the log-returns of $N$ assets represented by the random variables $X_{it}=\log\frac{P_{i,t+1}}{P_{i,t}}$, where $P_{i,t}$ is the price at time $t$, and the corresponding standardized variables $\tilde{X}_{it}$, with $i=1,...,N$ and $t=1,...,T$. The matrix of standardized returns $\mathbf{\tilde{X}}\in{\mathbb{R}}^{N\times T}$ is:
$\mathbf{\tilde{X}}=\left[\begin{array}[]{cccc}\tilde{X}_{11}&\tilde{X}_{12}&\dots&\tilde{X}_{1T}\\ \vdots&\vdots&\vdots&\vdots\\ \tilde{X}_{N1}&\tilde{X}_{N2}&\dots&\tilde{X}_{NT}\\ \end{array}\right].$ (1)
Since the rows of $\mathbf{\tilde{X}}$ are standardized, the correlation matrix is $\mathbf{C}=\frac{1}{T}\mathbf{\tilde{X}}\mathbf{\tilde{X}}^{T}.$
We denote by $(\lambda_{i},\bf{\phi}_{i})$, $i=1\dots N$, the eigenpairs of the matrix $\mathbf{C}$. Recall that $\sqrt{T\cdot\lambda_{i}}$ are the singular values of the matrix $\mathbf{\tilde{X}}$, and the number of nonzero singular values of $\mathbf{\tilde{X}}$ defines its rank. Since correlations are inherently signed, neglecting the diagonal elements – that are all equal to $1$ – the matrix $\bf C$ can be viewed as the weighted adjacency matrix of a signed network $G_{C}=(V,E,\mathbf{\mathbf{C}})$, where the nodes are the assets and the signed weighted edges are the correlations between them. We will call $G_{C}$ the correspondent correlation network.
@@ 3.1 A non-conservative information diffusion process on signed networks
We assume that the state of a node in the network is described by a variable that can take positive or negative values. A high absolute value of this variable reflects a high information content. A positive (negative) value on a node is interpreted as a favorable (unfavorable) information content for that node. An edge transfers its information between the two incident nodes. Negative edges flip the sign of walkers crossing them, while their sign remains unchanged when crossing positive edges. In other words, a positive walker becomes negative after crossing a negative edge, and is received by the destination node as the opposite of that in the source node, while it preserves its sign when crossing a positive edge. For instance, in a correlation network, if a piece of information is favorable to a given security, it remains favorable to all the positive correlated securities and it becomes unfavorable to the securities negatively correlated with the former. On the opposite, if a piece of information is unfavorable to a given security, it remains unfavorable to all the positively correlated securities, while it becomes favorable to the securities negatively correlated. An overall neutral information content on a node is represented by a state equal to zero. In general, although we refer to a discrete time process, the state variable on each node is assumed to be continuous.
@@
In the process described above, a node retains its own information and diffuses all its content to its neighbors. More interestingly, let us suppose that, at time $t$, a node retains all what it has received up to time $t-1$ from its neighbors and it does not transmit its entire information content but only the portion just received from its neighbors in the previous step $t-1$. In words, the node does not share all that its information content but only the part that it just learned from its neighbors. This is because what it knew before has already been sent in the previous steps of the diffusion process. Moreover, to ensure a higher generality, we assume that one node transmits only a fraction $\alpha(t)$ ($0<\alpha(t)\leq 1$) of what received. In summary, the dynamics is expressed by the following iterative relation, for $t\geq 0$,
@@
The above derivation provides an alternative interpretation of the exponential operator. In the literature (see [36]), the exponential operator usually measures the information diffusion in a Markovian context at $t=1$. We have shown that the same operator can also be interpreted as the equilibrium distribution of a non-Markovian process. This overcomes the issue of arbitrary tuning the time parameter at $t=1$.
@@ 3.2 Balance indices
where $[e^{\bf A}]_{i}$ denotes the $i-$th column of the exponential matrix. The amount of information originating from node $i$ that returns to the same node after the diffusion process and reaches a steady state is then given by the diagonal element
@@
This quantity is called local balance of the node $i$ and has been introduced in [36] by both a combinatoric perspective and a diffusive approach based on the Lerman-Gosh Laplacian [47]. In this context, it compares the fraction of information originating from $i$ that comes back to $i$ in the actual network and in the balanced versions of the same network. Equivalently, the numerator in Eq. (9) represents the polarization (see [46]) at each node, defined as the difference between the number of positive and negative walkers; the denominator is the total number of walkers coming back to that node. [^1] Similarly, the term $(e^{A})_{ij}$ captures the net influence that an initial shock in node $i$ exerts on node $j$. This object, known as signed communicability, has recently been proposed in [34] and could be used to define an edge balance index: $\kappa_{ij}:=\frac{(e^{A})_{ij}}{(e^{|A|})_{ij}}$. Such an index captures the effect of the signed edges and the unbalance on the transmission of information between nodes.
@@
The global balance index of the network is obtained by summing the contributions of all nodes:
@@
Finally, we observe that the above result still holds if we replace the binary matrix $\bf A$ with the weighted matrix $\bf W$. Moreover, the presence of identical loops in the network does not affect the computation of the balance indices. In other words, if the diagonal elements of the matrix are all equal, they do not contribute to the global balance, as shown in the following proposition.
Proposition 1.
Let $\bf A$ be a square matrix with real, possibly negative, entries and diagonal elements equal to $\chi\in{\mathbb{R}}$. Let $\tilde{\bf A}={\bf A}-\chi{\bf I}$, where ${\bf I}$ is the identity matrix, and $G$ and $\tilde{G}$ the corresponding graphs. Then $\kappa_{v}(G)=\kappa_{v}({\tilde{G}})$.
Proof.
@@ 4.1 Interpretation of the balance indices in terms of network structural reliability/predictability
These toy examples illustrate two important points. First, even within a fully deterministic diffusive model, the response of a node to state perturbations of at least two other nodes exhibits non-monotonic behaviors in an unbalanced network. Second, in a balanced network, there is a more pronounced dependence on the fluctuations in the attributes of some nodes than in an unbalanced network.
@@ 4.2 Global balance and condition number
In particular, we aim to discriminate when ${\mathscr{R}}({\bf A})$ is greater or less than $1$. In fact, if ${\mathscr{R}}({\bf A})<1$, the linear system in Eq. (6) on the signed network is better conditioned than the same system on the unsigned network, and vice versa if ${\mathscr{R}}({\bf A})>1$. This determines the different response of the nodes to perturbations propagating within the network in the two versions, signed and unsigned, of the network. We also show that ${\mathscr{R}}({\bf A})$ is tightly related to the global balance $\kappa(G)$ in Eq. (10) and, under some further assumptions, the time series of their values in a time-evolving network are highly correlated (see Sec. 5). We start with the following
Proposition 2.
The ratio ${\mathscr{R}}({\bf A})$ between the condition numbers of the matrices $e^{-{\bf A}}$ and $e^{-|{\bf A}|}$, in the Schatten $p-$norm with $p=1$, can be expressed as a function of the global balances of the networks $G$, $-G$ and $-|G|$ as
@@
where $-G$ is obtained by flipping the signs of all the edges in $G$, and $-|G|$ is the fully negative network.
Proof.
From Eqs. (10) and (15), we have
@@
Proposition 2 shows that the ratio of the two condition numbers is closely related to the global balance of the network $G$, also through the global balances of the networks associated with it, namely $-G$ and $-|G|$.
We now prove a necessary condition for $\kappa(G)=1$. First, observe that if $G$ is balanced, namely $\kappa(G)=1$, both $-G$ and $-|G|$ are antibalanced. Actually, we do not know whether they are balanced or not, that is, we can only say that $\kappa(-G)\leq 1$ and $\kappa(-|G|)\leq 1$. Nevertheless, the following result holds.
@@
Proof.
$=\frac{\left[\sum_{k}\frac{1}{2k!}\left(W^{+}_{2k}-W^{-}_{2k}\right)+\sum_{k}\frac{1}{(2k+1)!}\left(W^{+}_{2k+1}-W^{-}_{2k+1}\right)\right]\left[\sum_{k}\frac{1}{2k!}\left(W^{+}_{2k}-W^{-}_{2k}\right)-\sum_{k}\frac{1}{(2k+1)!}\left(W^{+}_{2k+1}-W^{-}_{2k+1}\right)\right]}{\left[\sum_{k}\frac{1}{2k!}\left(W^{+}_{2k}+W^{-}_{2k}\right)+\sum_{k}\frac{1}{(2k+1)!}\left(W^{+}_{2k+1}+W^{-}_{2k+1}\right)\right]\left[\sum_{k}\frac{1}{2k!}\left(W^{+}_{2k}+W^{-}_{2k}\right)-\sum_{k}\frac{1}{(2k+1)!}\left(W^{+}_{2k+1}+W^{-}_{2k+1}\right)\right]}$
where $W^{\pm}_{\rm even}$ and $W^{\pm}_{\rm odd}$ denote the total penalized number (weight) of positive and negative closed walks of even and odd length, respectively. For a balanced graph, by definition, there are no negative closed walks and $W^{-}_{\rm even}=W^{-}_{\rm odd}=0$. For an antibalanced graph, by definition, every even closed walk is positive and every odd closed walk is negative, so $W^{-}_{\rm even}=W^{+}_{\rm odd}=0$. The claim follows. ∎
We conjecture that Proposition 3 can also be extended to a sufficient condition, that is: If ${\mathscr{R}}({\bf A})=1$ then $G$ is a balanced or antibalanced signed network. [^2] We have numerically checked that all complete signed networks that can be generated up to a certain order $N$ satisfy this condition as both a necessary and a sufficient condition.
Observe that the previous Proposition expresses the ratio ${\mathscr{R}}({\bf A})$ as a function of the properties of the network $G$, namely through the penalized number or weight of even/odd and positive/negative closed walks in $G$. In the next Remark we highlight the explicit expression of these four types of closed walks.
@@
We now show that ${\mathscr{R}}({\bf A})$ is greater or less than $1$ according to the fact that the signed graph is even-dominant or odd-dominant.
Proposition 4.
Let ${\mathscr{R}}({\bf A})$ be the ratio between the condition numbers of the two matrices $e^{-{\bf A}}$ and $e^{-|{\bf A}|}$, in the Schatten $p-$norm with $p=1$. ${\mathscr{R}}({\bf A})<1$ if and only if the matrix ${\bf A}$ is the adjacency matrix of a strictly unbalanced even-dominant signed network $G$. ${\mathscr{R}}({\bf A})>1$ if and only if the matrix ${\bf A}$ is the adjacency matrix of a strictly unbalanced odd-dominant signed network $G$.
Proof.
We prove the claim for an even-dominant signed network. The odd-dominant case follows analogously. By expanding the ratio, we observe that:
${\mathscr{R}}({\bf A})=\frac{\left[W^{+}_{\rm even}-W^{-}_{\rm even}\right]^{2}-\left[W^{+}_{\rm odd}-W^{-}_{\rm odd}\right]^{2}}{\left[W^{+}_{\rm even}+W^{-}_{\rm even}\right]^{2}-\left[W^{+}_{\rm odd}+W^{-}_{\rm odd}\right]^{2}}<1\$
if and only if $W^{+}_{\rm even}\cdot W^{-}_{\rm even}-W^{+}_{\rm odd}\cdot W^{-}_{\rm odd}>0$. ∎
Proposition 2 shows the close relation between the ratio of the condition numbers in the signed and unsigned networks and the balance of the network $G$ and its related networks $-G$, $|G|$ and $-|G|$. Proposition 3 proves that, if the network $G$ is balanced or antibalanced, the conditioning of the linear problem in Eq. (6) is the same in the signed network and in its underlying unsigned network. Proposition 4 shows that this ratio can be both less or greater than $1$ depending on the topological properties of the network, in particular on the dominance of even closed walks or odd closed walks. Propositions 3 and 4 together provide an almost complete classification of the signed graphs based on the value of the ratio in Eq. (19). If ${\mathscr{R}}({\bf A})<1$ the graph is even-dominant strictly unbalanced; if ${\mathscr{R}}({\bf A})>1$ the graph is odd-dominant strictly unbalanced; if the graph is balanced or antibalanced, ${\mathscr{R}}({\bf A})=1$.

### v2

@@ Abstract
[…]
The global balance index is used in the network literature to quantify how balanced a signed network is. In this paper we show that the global balance index of financial correlation networks can be used as a systemic risk measure. We define the global balance index of a network starting from a diffusive process that describes how the information spreads across nodes in a network, providing an alternative derivation to the usual combinatorial one. The steady state of this process is the solution of a linear system governed by the exponential of the replication matrix of the process. We provide a bridge between the numerical stability of this linear system, measured by the condition number in an opportune norm, and the structural predictability of the underlying signed network. The link between the condition number and related systemic risk measures, such as the market rank indicators, allows the global balance index to be interpreted as a new systemic risk measure. A comprehensive empirical application to real financial data finally confirms that the global balance index of financial correlation networks represents a valuable and effective systemic risk indicator.
@@
Networks, Signed networks, Global balance, Systemic risk measures
@@ 1 Introduction
[…]
@@
[…]
In the literature on financial networks, correlation coefficients are often mapped onto the Euclidean space by introducing suitable distances. The most widely used methodology is based on the nonlinear transformation $d(i,j)=\sqrt{2(1-c_{ij})}$, introduced in [15], where $c_{ij}$ is the Pearson correlation between $i$ and $j$. Other transformations have been proposed to overcome some drawbacks of the previous one, such as the nonmetric distance $d(i,j)=-\log c_{ij}^{2}$ in [16]. While they both allow the construction of weighted undirected networks and the application of classical network theory tools, the former forbids indirect links to be shorter than direct links, and the latter fails to exploit the information contained in the correlation sign.
@@
The theory of signed networks [37] and their balance can provide interesting insight into the subtle relationship between predictability and overall systemic risk associated with groups of stocks or the entire network. A key insight, due to Harary, is that the level of structural balance impacts also on the performance of an investment portfolio. In fact, to the best of our knowledge, the first paper that establishes a connection between signed networks and risk in financial networks is the one by Harary et al. [38]. In this paper, the authors address the question of how to design an effective hedging strategy for portfolio selection. They analyze the network of correlations between securities and conclude that a hedging strategy is effective, and thus the portfolio less risky, if there is at least one negative link and if the signed network is structurally balanced. However, the authors do not deal with systemic risk measures in the proper sense and do not study the relationship between these measures and structural balance indices. Indeed, the global balance index has never been interpreted as an indicator of systemic risk.
@@
In this research, we show how to move from these structural indicators for signed networks to systemic risk measures as usually interpreted in the literature. The logical path of the paper can be described as follows. We start from a non-conservative diffusive process on networks governed by the adjacency matrix, interpreted as a replication matrix of the information transmitted by the nodes. We interpret the local version of the structural balance as a measure of the amount of information that leaves one node and returns to the same node when the diffusion process has reached the steady state. From the local measure we move to the global balance index, which is known and widely used in the literature to quantify structural balance in signed networks. Then, we show that this global indicator is naturally related to a ratio of the condition numbers of the two linear systems defining the steady state of the diffusive process on the signed network and on the corresponding unsigned network, that is the network with all positive edges. Thus, a natural bridge is created between the structural balance in correlation networks and a measure of the sensitivity of the network to an external perturbation of the state of the nodes. The link to known algebraic measures of conditioning provides the further connection to systemic risk measures. The condition number of the correlation matrix and some suitable generalizations called market rank indicators (MRI) have already been proposed in the literature as systemic risk measures (see [39]). We then compare the global balance of the signed correlation network with these systemic risk measures, which are part of the broader class of so-called proper measures of connectedness. This comparison leads us to explore the ability of the global balance to discriminate systemic events. Finally, we look at how such a global indicator performs as an effective measure of systemic risk, testing its performance on real financial data.
@@
The paper is structured as follows. In Section 2 we recall the mathematical definitions. In Section 3 we show how the structural balance can be obtained from a discrete diffusive process on the network. In Section 4 we outline a conceptual path from the balance indices to systemic risk measures. In Section 5 we perform an empirical analysis to prove the effectiveness of the balance index as systemic risk measure. In Section 6 we draw conclusions and outlines for future research.
@@ 2.1 Signed networks and structural balance
[…]
A weighted undirected network is represented by a graph $G=(V,E,\mathbf{W})$, where $V$ is the set of $N$ nodes, and $E$ is the set of undirected edges. The weighted adjacency matrix $\mathbf{W}=[w_{ij}]$ of order $N\times N$, with $w_{ij}\in\mathbb{R}^{+}$, encodes all the adjacency relationships between nodes. Specifically, $w_{ij}>0$ if there is an edge between nodes $i$ and $j$, $w_{ij}=0$ otherwise. A weighted undirected network is said to be signed if the edge weights can be positive or negative, that is $w_{ij}\in\mathbb{R}$. The edge signs depend on the nature of the relationship between the pair of nodes. It may relate to activating or inhibiting functions in biological systems [40], trusting or distrusting links in social or political networks [41], cooperative or antagonistic relationships in economic settings [42], correlated and anticorrelated behavior in the time series of asset returns in the financial context [43].
A binary adjacency matrix $\mathbf{A}=[a_{ij}]$ can be associated with the signed network $G$ by setting all non-null weights equal to $1$ or $-1$ according to the sign of the entries $w_{ij}$. In the paper we will also need to refer to the underlying unsigned network, that is the graph obtained from $G$ by neglecting the edges sign. The matrices associated with this network are $|{\bf W}|$ and $|{\bf A}|$. The strength of a node $i$, in the signed network, is then defined as $s_{i}=\sum_{j}|w_{ij}|$ and the degree as $d_{i}=\sum_{j}|a_{ij}|$.
A signed graph $G$ is structurally balanced if there are no negative cycles, that is, cycles with an odd number of negative edges (see [17, 18]). A balanced network is characterized by the balance theorem as a graph that admits a bipartition of the vertex set $V$ such that every edge between the two subsets is negative, while every edge within each subset is positive (see [17]).
@@ 2.2 Correlation networks from asset returns
Let us consider the time series of the log-returns of $N$ assets represented by the random variables $X_{it}=\log\frac{P_{i,t+1}}{P_{i,t}}$, where $P_{i,t}$ is the price at time $t$, and the corresponding standardized variables $\tilde{X}_{it}$, with $i=1,...,N$ and $t=1,...,T$. Given the matrix of the standardized returns $\mathbf{\tilde{X}}\in{\mathbb{R}}^{N\times T}$, the correlation matrix is $\mathbf{C}=\frac{1}{T}\mathbf{\tilde{X}}\mathbf{\tilde{X}}^{T}.$ We denote by $(\lambda_{i},{\bf{\phi}}_{i})$, $i=1\dots N$, the eigenpairs of the matrix $\mathbf{C}$. Recall that $\sqrt{T\cdot\lambda_{i}}$ are the singular values of the matrix $\mathbf{\tilde{X}}$, and the number of nonzero singular values of $\mathbf{\tilde{X}}$ defines its rank. Since correlations are inherently signed, neglecting the diagonal elements – that are all equal to $1$ – the matrix $\bf C$ can be viewed as the weighted adjacency matrix of a signed network $G_{C}=(V,E,\mathbf{\mathbf{C}})$, where the nodes are the assets and the signed weighted edges are the correlations between them. We will call $G_{C}$ the corresponding correlation network.
@@ 3.1 A non-conservative information diffusion process on signed networks
We assume that the state of a node in the network is described by a variable that can take positive or negative values. A high absolute value of this variable reflects a high information content. A positive (negative) value on a node is interpreted as a favorable (unfavorable) information content for that node. An edge transfers information between the two incident nodes. Negative edges flip the sign of walkers crossing them, while their sign remains unchanged when crossing positive edges. In other words, a positive walker becomes negative after crossing a negative edge, and is received by the destination node as the opposite of that in the source node, while it preserves its sign when crossing a positive edge. For instance, in a correlation network, if a piece of information is favorable to a given security, it remains favorable to all the positive correlated securities and it becomes unfavorable to the securities negatively correlated with the former. On the opposite, if a piece of information is unfavorable to a given security, it remains unfavorable to all the positively correlated securities, while it becomes favorable to the securities negatively correlated. An overall neutral information content on a node is represented by a state equal to zero. In general, although we refer to a discrete time process, the state variable on each node is assumed to be continuous.
@@
[…]
@@
[…]
@@ 3.2 Balance indices
where $[e^{\bf A}]_{i}$ denotes the $i-$th column of the exponential matrix. The amount of information originating from node $i$ that returns to the same node after the diffusion process reaches a steady state is then given by the diagonal element
@@
This quantity, called local balance of the node $i$, has been introduced in [36] by both a combinatoric perspective and a diffusive approach based on the Lerman-Gosh Laplacian [47]. In this context, it compares the fraction of information originating from $i$ that comes back to $i$ in the actual network and in the balanced versions of the same network. The numerator in Eq. (8) represents the polarization (see [46]) at each node, defined as the difference between the number of positive and negative walkers; the denominator is the total number of walkers coming back to that node. [^1] Similarly, the term $(e^{A})_{ij}$ captures the net influence that an initial shock in node $i$ exerts on node $j$. This object, known as signed communicability, has recently been proposed in [34] and could be used to define an edge balance index: $\kappa_{ij}:=\frac{(e^{A})_{ij}}{(e^{|A|})_{ij}}$. Such an index captures the effect of the signed edges and the unbalance on the transmission of information between nodes.
@@
The global balance index of the signed network $G$ is defined as:
@@
Finally, we observe that the above result still holds if we replace the binary matrix $\bf A$ with the weighted matrix $\bf W$. Moreover, the presence of identical loops in the network does not affect the computation of the balance indices. In other words, if the diagonal elements of the matrix are all equal, they do not contribute to the global balance.
More specifically, let $\bf A$ be a square matrix with real, possibly negative, entries and diagonal elements equal to $\chi\in{\mathbb{R}}^{+}$. Let $\tilde{\bf A}={\bf A}-\chi{\bf I}$, where ${\bf I}$ is the identity matrix, and $G$ and $\tilde{G}$ the corresponding graphs. Then $\kappa_{v}(G)=\kappa_{v}({\tilde{G}})$. In fact,
@@ 4.1 Interpretation of the balance indices in terms of network structural reliability/predictability
These examples illustrate two important points. First, even within a fully deterministic diffusive model, the response of a node to state perturbations of at least two other nodes exhibits non-monotonic behaviors in an unbalanced network. Second, in a balanced network, there is a more pronounced dependence on the fluctuations in the attributes of some nodes than in an unbalanced network.
@@ 4.2 Global balance and condition number
In particular, we aim to discriminate when ${\mathscr{R}}({\bf A})$ is greater or less than $1$. In fact, if ${\mathscr{R}}({\bf A})<1$, the linear system in Eq. (5) on the signed network is better conditioned than the same system on the unsigned network, and vice versa if ${\mathscr{R}}({\bf A})>1$. This determines the different response of the nodes to perturbations propagating within the network in the two versions, signed and unsigned, of the network. We also show that ${\mathscr{R}}({\bf A})$ is tightly related to the global balance $\kappa(G)$ in Eq. (9) and, under some further assumptions, the time series of their values in a time-evolving network are highly correlated (see Sec. 5). We start observing that the ratio ${\mathscr{R}}({\bf A})$ between the condition numbers of the matrices $e^{-{\bf A}}$ and $e^{-|{\bf A}|}$, in the Schatten $p-$norm with $p=1$, can be expressed as a function of the global balances of the networks $G$, $-G$ and $-|G|$ as
@@
where $-G$ is obtained by flipping the signs of all the edges in $G$, and $-|G|$ is the fully negative network. In fact, from Eqs. (9) and (14), we have
@@
Formula 16 shows that the ratio of the two condition numbers is closely related to the global balance of the network $G$, also through the global balances of the networks associated with it, namely $-G$ and $-|G|$.
[…]
@@
The proof of this proposition can be found in the Supplementary Material, Sec. 2. We conjecture that this necessary condition can also be extended to a sufficient condition, that is: If ${\mathscr{R}}({\bf A})=1$ then $G$ is a balanced or antibalanced signed network. [^2] We have numerically checked that all complete signed networks that can be generated up to a certain order $N$ satisfy this condition as both a necessary and a sufficient condition.
[…]
${\mathscr{R}}({\bf A})=\frac{\left[W^{+}_{\rm even}-W^{-}_{\rm even}\right]^{2}-\left[W^{+}_{\rm odd}-W^{-}_{\rm odd}\right]^{2}}{\left[W^{+}_{\rm even}+W^{-}_{\rm even}\right]^{2}-\left[W^{+}_{\rm odd}+W^{-}_{\rm odd}\right]^{2}}$ (17)
[…]
where $W^{\pm}_{\rm even}$ and $W^{\pm}_{\rm odd}$ denote the total penalized weight of positive and negative closed walks of even and odd length, respectively.
In the next Remark we highlight the explicit expression of these four types of closed walks.
@@
[…]
where ${\lambda}_{N}\leq{\lambda}_{N-1}\leq\dots\leq{\lambda}_{1}$ and $\overline{\lambda}_{N}\leq\overline{\lambda}_{N-1}\leq\dots\leq\overline{\lambda}_{1}$ are the eigenvalues of $\bf A$ and $\bf|A|$, respectively.
@@
We can now show that ${\mathscr{R}}({\bf A})$ is greater or less than $1$ according to the fact that the signed graph is even-dominant or odd-dominant. Let ${\mathscr{R}}({\bf A})$ be the ratio between the condition numbers of the two matrices $e^{-{\bf A}}$ and $e^{-|{\bf A}|}$, in the Schatten $p-$norm with $p=1$. ${\mathscr{R}}({\bf A})<1$ if and only if the matrix ${\bf A}$ is the adjacency matrix of a strictly unbalanced even-dominant signed network $G$. ${\mathscr{R}}({\bf A})>1$ if and only if the matrix ${\bf A}$ is the adjacency matrix of a strictly unbalanced odd-dominant signed network $G$. In fact, if we refer for instance to an even-dominant signed network, by expanding the ratio as in Eq. 17, we have that ${\mathscr{R}}({\bf A})<1$ if and only if $W^{+}_{\rm even}\cdot W^{-}_{\rm even}-W^{+}_{\rm odd}\cdot W^{-}_{\rm odd}>0$.
Formula 16 shows the close relation between the ratio of the condition numbers in the signed and unsigned networks and the balance of the network $G$ and its related networks $-G$, $|G|$ and $-|G|$. If the network $G$ is balanced or antibalanced, the conditioning of the linear problem in Eq. (5) is the same in the signed network and in its underlying unsigned network. The ratio ${\mathscr{R}}({\bf A})$ can be both less or greater than $1$ depending on the topological properties of the network, in particular on the dominance of even closed walks or odd closed walks. The previous considerations provide an almost complete classification of the signed graphs based on the value of the ratio in formula (17). If ${\mathscr{R}}({\bf A})<1$ the graph is even-dominant strictly unbalanced; if ${\mathscr{R}}({\bf A})>1$ the graph is odd-dominant strictly unbalanced; if the graph is balanced or antibalanced, ${\mathscr{R}}({\bf A})=1$.